what is the difference?
-5^2 - I thought the answer is 25
(-5)^2 - I thought the answer is 25
-(5)^2 - I thought the answer is -25

Answer by checkley77(12844)

-5^2 - I thought the answer is 25
(-5)^2 - I thought the answer is 25
-(5)^2 - I thought the answer is -25
--------------------
-5^2=-(5*5)=-25 ans.
(-5)^2=(-5*-5)=25 ans.
-(5)^2=-(5*5)=-25 ans.